# Query-Specific GNN: A Comprehensive Graph Representation Learning Method for Retrieval Augmented Generation

### v1

@@ A.8 Good Case Studies
We present three typical comparative cases between QSGNN and HippoRAG2 to demonstrate QSGNN’s advantages. In each case, correct retrieval evidence is marked in green and incorrect evidence in red. Notably, QSGNN’s marks are based on query alignment, whereas HippoRAG2’s marks is based on PPR scores  (Gutiérrez et al., 2025).
Figure 4: Good case 1 for QSGNN.
Figure 5: Good case 2 for QSGNN.
Figure 6: Good case 3 for QSGNN.
Good Case 1. Figure  4 illustrates a scenario where QSGNN achieves recall@5 of 1.0 with correct answer generation, while HippoRAG2 fails. The 4-hop question requires understanding composite semantics: i) “the only group larger than” corresponds to “the world’s second largest recorded music company”, ii) “the headquarters of” refers to “based in Santa Monica, California” and iii) “the explorer reach the city” denotes “set foot in the area was the party of explorer Gaspar de Portola”. QSGNN successfully identifies the critical documents describing Sony Music Entertainment’s ranking, location and the city information, whereas HippoRAG2 erroneously retrieves information about Jive Records’ headquarters. This failure stems from HippoRAG2’s selection of seed nodes (“headquarter”, “located in”) without high-level contextual information. The case validates Multi-L KG’s semantic modeling capability and query alignment’s effectiveness in filtering misinformation.
Good Case 2. Figure  5 shows QSGNN achieving perfect recall@5 while retrieval of HippoRAG2 is inaccurate. For the 4-hop question about Andrew Deveaux’s birthplace, QSGNN integrates evidence across multiple documents and each document describes one key evidence for the question: i) biographical data (“born in South Carolina”), ii) state capital history (“Columbia became capital in 1786”), and iii) geographic relationships (“Forest Acres borders Columbia”). The successful retrieval is due to the query-alignment and comprehensive understanding of contextual information. However, HippoRAG2 incorrectly retrieves Savannah county information due to over-reliance on “county” and “located on” seed nodes.
Good Case 3. Figure  6 presents a good case of QSGNN where it retrieves correctly and gives the correct answer but HippoRAG2 can’t answer correctly because of the inaccurate retrieval. In this case, QSGNN successfully find the snippets for the “where the Yongle Emperor greeted the Karmapa”, “the capitol city of Yaxing Coach’s headquarters location” and the history information about “Yangzhou”, which is the combination results of Multi-L KG and query-specific aggregation. HippoRAG2 retrieves redundant Sino-Tibetan relations documents. The failure occurs because HippoRAG2 prioritizes high PPR scores for entities “Yongle Emperor” and “Karmapa” without considering their contextual relationships.
A.9 Bad Case Discussion and Limitations
We analyze three representative failure cases of QSGNN in document retrieval, each illustrating distinct limitations that inform our future research directions.
Figure 7: Bad case 1 for QSGNN.
Figure 8: Bad case 2 for QSGNN.
Figure 9: Bad case 3 for QSGNN.
@@ A.9 Bad Case Discussion and Limitations
Bad Case 1. Figure  7 demonstrates QSGNN’s failure to retrieve documents containing “Freikorps” due to a misspelled query terminology (“free crops” should be “Freikorps”, we changed the “free crops” to “Freikorps” then QSGNN could retrieve correctly). This spelling error caused QSGNN to prioritize non-critical entities like “democratic government” and “Germany”. We also check the pre-training data and fine-tuning data, there is no information about the associations between “free crops” and “Freikorps”, which exacerbates the model’s confusion.
Bad Case 2. Figure  8 reveals QSGNN’s limitation in processing specific term (“San Clemente”, “Fuser”, “Alberto”), which may not be well represented by QSGNN since they are absent from both pre-training and fine-tuning corpora. The model deflect to general terms like “province” and “county”, retrieving irrelevant documents.
Bad Case 3. Figure  9 shows a typical bad case for QSGNN. In this case the term like “Messi” and “Barcelona” dwarf the key evidence “Diego Maradona, who Messi can bring comparison to”. The sequential dependency between finding “Diego Maradona” and subsequent evidence (“June 1982 transfer record”) further complicates retrieval. It is difficult for query alignment to identify the correct information because huge amount of documents describing “Messi” and “Barcelona”, which is also related to query, overwhelm the key snippets of “Diego Maradona”.
Limitation Discussion and Future Work. Based on the analysis of the bad cases above, QSGNN is not a silver bullet for all multi-hop scenarios. It exhibits limitations when dealing with specific terminologies especially when they are not well presented in the corpora (see bad case 1, 2), and also struggles to deal the situation where the prevalent mis-related documents overwhelm the key evidence with sequential answer dependencies (see bad case 3). We think these limitations can be mitigated by two strategies: i) For the first issue, selecting seed nodes as domain-specific terminologies and sampling constrained subgraphs(depending on the layer of QSGNN, in our case the longest hop is 4), could prioritize relevant documents in retrieval, though this risks information loss. ii) For the second issue, integrating Chain-of-Thought(CoT) reasoning  (Chen et al., 2019; Lyu et al., 2023) through query decomposition and iterative retrieval may mitigate the problem. However, the first solution is a tricky strategy and even may compromise the performance since subgraph sampling will lead to information loss. The second solution introduces CoT into framework, which is beyond the design topic of QSGNN. We will leave the combination of these two possible solutions without compromising the performance as the future work of QSGNN.
A.10 Prompts for OpenIE
We present the prompts we used for OpenIE workflow. Figure  10 shows the prompts used for sentence extraction, Figure  11 shows the prompts for entity extraction, and Figure  12 shows the prompts for triple extraction. First, we extract sentences from each document. Then we identify core entities from extracted sentences. At last, these entities are used to formulate triples from the document.
Figure 10: Sentence extraction prompt for OpenIE.
Figure 11: Entity extraction prompt for OpenIE.
Figure 12: Triple extraction prompt for OpenIE.
A.11 Prompts for Synthesized Pre-training Data
The prompts used for generate pre-training data are shown in Figure  13 and Figure  14.
Figure 13: One hop question generation prompt for OpenIE.
Figure 14: Two hop question generation prompt for OpenIE.

### v2

@@ A.4. Case Study and Discussion
We present the case study of Quest-GNN here, which includes two good cases and two bad cases. Correct evidence is marked in green and incorrect evidence in red. Notably, Quest-GNN’s marks are based on question alignment, whereas HippoRAG2’s marks are based on PPR scores  (Gutiérrez et al., 2025).
Figure 4. Good case 1 for Quest-GNN.
Figure 5. Good case 2 for Quest-GNN.
Good Case Analysis. Figure  4 and Figure  5 show two good cases where Quest-GNN achieves recall@5 of 1.0 with correct answer generation while retrieval of HippoRAG2 is inaccurate. Quest-GNN successfully identifies the critical understanding composite semantics (For example, “the headquarters of” refers to “based in Santa Monica, California” in case 1), and integrates evidence across multiple documents (For example, biographical data:“born in South Carolina”, state capital history:“Columbia became capital in 1786”, and geographic relationships:“Forest Acres borders Columbia” in case 2). But HippoRAG2 fails to select the right seed nodes without high-level contextual information. These two cases show Multi-L KG’s semantic modeling capability and question alignment’s effectiveness in filtering misinformation and understanding of contextual information comprehensively.
Figure 6. Bad case 1 for Quest-GNN.
Figure 7. Bad case 2 for Quest-GNN.
@@
Bad Case Analysis. Figure  6 and Figure  7 reveal two bad cases for Quest-GNN. These two cases both stem from poor initial representation. In case 1 “Freikorps” is a misspelled query terminology, which should be “free crops” (we changed to the “free crops” then Quest-GNN could retrieve correctly). This leads Quest-GNN to prioritize non-critical entities like “democratic government” and “Germany”. We also check the pre-training data and fine-tuning data, there is no information about the associations between “free crops” and “Freikorps”, which exacerbates the model’s confusion. In case 2, “San Clemente”, “Fuser” and “Alberto” are all under-represented since they are absent from both pre-training and fine-tuning corpora. However, all of them are important for correct retrieval.
Limitation Discussion and Future Work. Based on the analysis of the bad cases above, Quest-GNN exhibits limitations when dealing with specific terminology or misspelling that can not be well-represented. We think this limitation can be mitigated by selecting seed nodes as domain-specific terminologies and sampling constrained subgraphs to prioritize relevant documents in retrieval. But how to realize this with limited information loss is still non-trivial. We will leave this as our future work.
A.5. Prompts
The prompts for OpenIE and question generation are in Figure  8.
Figure 8. Extraction and question generation prompts.